OpenStudy (anonymous):

how does (x^3-8)=(x-2)(x^2+2x+4)??

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{it is the famous formula}\\~\\ \rm(a^3-b^3)=(a-b)(a^2+ab+b^2)}\end{align}\)

OpenStudy (mathmath333):

\(\large\rm\begin{align} \color{black}{x^3-2^3=(x-2)(x^2+2x+2^2)}\end{align}\)

OpenStudy (anonymous):

whats the formula called?